# Random Sperner lemma and random Brouwer fixed point theorem

### v2

@@ 1 Introduction
Throughout, $\mathbb{R}^{d}$ denotes the $d$-dimensional Euclidean space. The well known Brouwer fixed point theorem [3] states that every continuous mapping from a bounded closed convex subset of $\mathbb{R}^{d}$ to itself has a fixed point. Brouwer fixed point theorem has rich applications in analysis, geometry, differential equations and mathematical economics [37, 7, 12, 38], in particular, it is also attractive that Brouwer fixed point theorem is equivalent to many famous fundamental results [36, 27, 28, 39], for example, the Knaster-Kuratowski-Mazurkiewicz (briefly, $KKM$) mapping principle [31], the classical Borsuk theorem [2] and the Gustafson-Schmitt-Cid-Mawhin’s theorem on the existence of periodic solutions of ordinary differential equations [7, Theorem 1.1]. Due to its fundamental importance, Brouwer fixed point theorem has been generalized in various ways [30, 44, 48] and has had a major impact on the development of nonlinear analysis and fixed point theory [1, 8, 16].
@@
It is well known that the proof of the classical Brouwer fixed point theorem is not very easy and the simplest of all the classical proofs is by means of Sperner lemma [1, 31, 33, 46]. The classical Sperner lemma states that for any simplicial subdivision $\mathscr{S}$ of a simplex and any proper labeling function of $\mathscr{S}$ there are an odd number of completely labeled simplexes in $\mathscr{S}$. It is also well known that the most accessible proof of Sperner lemma was given by Kuhn in [33] via the graph-theoretic method. However, the proof of random Brouwer fixed point theorem involves more since $L^{0}(\mathcal{F},\mathbb{R}^{d})$ admits a more complicated topological and algebraic structure than $\mathbb{R}^{d}$, which requires the following three nontrivial tools. The first tool is Lemma 4.3 of [24], which states that a $\sigma$-stable mapping is $\mathcal{T}_{c}$-continuous iff it is random sequentially continuous. Since the topology $\mathcal{T}_{c}$ is too strong for us to conveniently treat the fixed point problem of the $\mathcal{T}_{c}$-continuous mappings, but random sequential continuity can be described in terms of $\mathcal{T}_{\varepsilon,\lambda}$, this lemma enables us to make full use of $\mathcal{T}_{\varepsilon,\lambda}$ to treat random Brouwer fixed point theorem. The second tool is the randomized version of the classical Bolzano-Weierstrass theorem [29], which states that an almost surely (briefly, a.s.) bounded sequence in $L^{0}(\mathcal{F},\mathbb{R}^{d})$ always admits an a.s. convergent randomized subsequence although it may admit no a.s. convergent subsequence. Since an a.s. bounded, $\mathcal{T}_{\varepsilon,\lambda}$-closed and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$, is generally noncompact but random sequentially compact by the randomized Bolzano-Weierstrass theorem, then the randomized Bolzano-Weierstrass theorem can be used to overcome the difficulty due to noncompactness [10, 24, 26]. The third tool is a random version of the classical Sperner lemma. To prove a special case of the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Brouwer fixed point theorem (see [10, Proposition 3.1]), the authors [10] first give a special case of random Sperner lemma (see the first part of [10, Theorem 2.3]), where the authors studied a special $L^{0}$-simplicial subdivision of an $L^{0}$-simplex (called the $L^{0}$-barycentric subdivision) in detail. The special random Sperner lemma provided in [10] was employed in all the works [10, 24, 26] on random Brouwer fixed point theorem. However, the proof of [10] on the special random Sperner lemma is not easily accessible to the reader, for example, the idea of the proof in [10] was very similar to Kuhn’s one [33], but the method in [10] was not by means of the graph-theoretic argumentation and thus rather complicated. Considering the fundamental importance of random Brouwer fixed point theorem, for example, it has been used in [24] for the proof of a noncompact Schauder fixed point theorem in random normed modules, and it will also be used to establish a noncompact $KKM$ principle in random locally convex modules so that a series of open problems posed in [24] will be solved in our forthcoming work, it is necessary to look for a very easily accessible proof of random Sperner lemma and to further give a complete proof of random Brouwer fixed point theorem. This leads us directly to systematically developing the theory of $L^{0}$-simplicial subdivisions of $L^{0}$-simplexes.
@@ 2.1 Different forms and special cases of random Brouwer fixed point theorem
It is obvious that a $\mathcal{T}_{\varepsilon,\lambda}$-continuous mapping defined on a $\sigma$-stable set is random sequentially continuous, Proposition 2.3 below, which first occurred in [25, Proposition 4.2] is a special case of Proposition 2.2.
@@
We would also like to point out that Proposition 2.2 (namely Theorem 2.1) and Proposition 2.3 are both best possible, which is explained as follows. It is well known that $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is a metrizable linear topological space under $\mathcal{T}_{\varepsilon,\lambda}$ (namely the topology of convergence in probability), there is, naturally, the notion of boundedness in the sense of the linear topology $\mathcal{T}_{\varepsilon,\lambda}$, denoted by the $\mathcal{T}_{\varepsilon,\lambda}$-boundedness, it is easy to see that the a.s. boundedness implies the $\mathcal{T}_{\varepsilon,\lambda}$-boundedness, and we have known from Proposition 16 of [22] that these two kinds of boundedness coincide for a $\sigma$-stable subset. Likewise, the $L^{0}$-convexity implies the usual convexity, and we have also known from Lemma 2.8 of [47] that these two kinds of convexity coincide for a $\mathcal{T}_{\varepsilon,\lambda}$-closed and $\sigma$-stable subset. On the one hand, an a.s. bounded $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset $G$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is generally noncompact, if the hypothesis on the $\sigma$-stability of $f$ in Proposition 2.3 is removed, then Proposition 2.3 does not hold, which comes from a classical result due to Dobrowolski and Marciszewski [9] (see also [24, Section 2.1]). On the other hand, an example provided by Ponosov in [42] shows that Proposition 2.4 (namely Proposition 2.3) does not hold if $G_{U}$ is replaced with a generic $\mathcal{T}_{\varepsilon,\lambda}$-closed, $\mathcal{T}_{\varepsilon,\lambda}$-bounded and convex set $G$ (namely $G$ is not $L^{0}$-convex). Therefore, they are all best possible.
@@ 2.2 A general $L^{0}$-simplicial subdivision and random Sperner lemma
In [10, Definition 2.1], Drapeau et al. introduced the notion of a proper $L^{0}$-labeling function for the $L^{0}$-barycentric subdivision of an $L^{0}$-simplex in $L^{0}(\mathcal{F},\mathbb{R}^{d})$. Definition 2.7 below is a generalization of [10, Definition 2.1]. But we adopt a strengthened notion of a completely labeled $L^{0}$-simplex since such a notion is both easy to understand and such a completely labeled $L^{0}$-simplex indeed exists (see Theorem 2.9). In fact, according to Definition 2.7, $\phi(y_{j})$ is an equivalence class of a constant valued random variable for each $j\in\{1,2,\cdots,n\}$ for a completely labeled $L^{0}$-simplex $G=Conv_{L^{0}}(\{y_{1},y_{2},\cdots,y_{n}\})$, and hence our notion of a completely labeled $L^{0}$-simplex is completely similar to the notion of a completely labeled simplex.
@@
We first give the following two Examples — Examples 3.1 and 3.2. The first one shows the $L^{0}$-barycentric subdivision introduced in [10, Definition 1.10] is a special case of our Definition 2.6. The second one shows our Definition 2.6 does not just include the $L^{0}$-barycentric subdivision.
@@
Then, by applying the same method as in [10, Lemma 1.11], one can verify that $\{S_{\pi}:\pi\in G(n)\}$ is an $L^{0}$-simplicial subdivision of $S$, which is called the $L^{0}$-barycentric subdivision of $S$.
@@
Definition 4.1 here is adopted from [17, 18] by following the traditional nomenclature of random metric spaces and random normed spaces (see [45, Chapters 9 and 15]).
@@
(1) In Definition 4.3, we require $n_{k}$ to be an element in $L^{0}(\mathcal{F},\mathbb{N})$, rather than a positive integer-valued random variable as in [24, Definition 2.10]. This choice is made for simplifying written expressions, and it is easy to check that the two formulations are essentially equivalent. (2) In the sense of Definition 4.3, the randomized Bolzano-Weierstrass theorem can be restated as: any a.s. bounded nonempty subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is random relatively sequentially compact. Moreover, as pointed out in [24], if $G$ is $\sigma$-stable, then $G$ is random relatively sequentially compact iff $\bar{G}_{\varepsilon,\lambda}$ (the closure of $G$ under $\mathcal{T}_{\varepsilon,\lambda}$) is random sequentially compact. Consequently, any a.s. bounded $\sigma$-stable nonempty $\mathcal{T}_{\varepsilon,\lambda}$-closed subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ (for example, an $L^{0}$-simplex) must be random sequentially compact. (3) For any $(n-1)$-dimensional $L^{0}$-simplex $S$ of an $RN$ module, as pointed out by [24, Lemma 5.5], there exists an $L^{0}$-affine $\mathcal{T}_{\varepsilon,\lambda}$-isomorphism $T$ from $\Delta_{n-1}$ to $S$, where $\Delta_{n-1}:=Conv_{L^{0}}(\{e_{1},e_{2},\cdots,e_{n}\})$ is the standard $(n-1)$-dimensional $L^{0}$-simplex of $L^{0}(\mathcal{F},\mathbb{R}^{n})$. Since $\Delta_{n-1}$ is random sequentially compact as discussed above, and $T$ is $\sigma$-stable and $\mathcal{T}_{\varepsilon,\lambda}$-continuous, then $S$ is random sequentially compact.
@@
It is clear that a.s. sequential continuity implies $\mathcal{T}_{\varepsilon,\lambda}$-continuity. Moreover, when $G_{1}$ is $\sigma$-stable, the $\mathcal{T}_{\varepsilon,\lambda}$-continuity of a mapping $T$ from $G_{1}$ to $G_{2}$ necessarily implies random sequential continuity of $T$. Furthermore, [24, Lemma 4.3] shows that a $\sigma$-stable mapping $T$ of $G_{1}$ to $G_{2}$ is random sequentially continuous iff it is $\mathcal{T}_{c}$-continuous.
@@
Lemma 4.9 is a generalization of [10, Theorem 2.3], since $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is an $RN$ module, and a $\sigma$-stable a.s. sequentially continuous mapping is, obviously, random sequentially continuous, as noted in Remark 4.6. Since random sequential continuity is much more complicated than a.s. sequential continuity, Lemma 4.4 of [24] has played a key role in the treatment of random sequential continuity in the proof of Lemma 4.9. Besides, we would also like to emphasize another difference between the proofs of Lemma 4.9 and [10, Theorem 2.3]: the sequence $\{S^{m},m\in\mathbb{N}\}$ of $(n-1)$-dimensional $L^{0}$-simplexes, which is constructed in Lemma 4.9, has the nice property that $S^{m+1}\subseteq S^{m}$ for each $m\in\mathbb{N}$, but the sequence $\{S^{m},m\in\mathbb{N}\}$ constructed in [10, Theorem 2.3] does not have this property. The nice property of our sequence will be used to study other problems in the future. Besides, the proof of Lemma 4.9 is simpler and clearer than that of [10, Theorem 2.3] since we only employ a single proper $L^{0}$-labeling function throughout the proof of Lemma 4.9.

### v3

@@ 1 Introduction
Throughout, $\mathbb{R}^{d}$ denotes the $d$-dimensional Euclidean space. The well known Brouwer fixed point theorem [3] states that every continuous mapping from a bounded closed convex subset of $\mathbb{R}^{d}$ to itself has a fixed point. Brouwer fixed point theorem has rich applications in analysis, geometry, differential equations and mathematical economics [39, 8, 14, 40], in particular, it is also attractive that Brouwer fixed point theorem is equivalent to many famous fundamental results [38, 29, 30, 41], for example, the Knaster-Kuratowski-Mazurkiewicz (briefly, $KKM$) mapping principle [33], the classical Borsuk theorem [2] and the Gustafson-Schmitt-Cid-Mawhin’s theorem on the existence of periodic solutions of ordinary differential equations [8, Theorem 1.1]. Due to its fundamental importance, Brouwer fixed point theorem has been generalized in various ways [32, 46, 50, 7, 13] and has had a major impact on the development of nonlinear analysis and fixed point theory [1, 9, 18].
@@
It is well known that the proof of the classical Brouwer fixed point theorem is not very easy and the simplest of all the classical proofs is by means of Sperner lemma [1, 33, 35, 48]. The classical Sperner lemma states that for any simplicial subdivision $\mathscr{S}$ of a simplex and any proper labeling function of $\mathscr{S}$ there are an odd number of completely labeled simplexes in $\mathscr{S}$. It is also well known that the most accessible proof of Sperner lemma was given by Kuhn in [35] via the graph-theoretic method. However, the proof of random Brouwer fixed point theorem involves more since $L^{0}(\mathcal{F},\mathbb{R}^{d})$ admits a more complicated topological and algebraic structure than $\mathbb{R}^{d}$, which requires the following three nontrivial tools. The first tool is Lemma 4.3 of [26], which states that a $\sigma$-stable mapping is $\mathcal{T}_{c}$-continuous iff it is random sequentially continuous. Since the topology $\mathcal{T}_{c}$ is too strong for us to conveniently treat the fixed point problem of the $\mathcal{T}_{c}$-continuous mappings, but random sequential continuity can be described in terms of $\mathcal{T}_{\varepsilon,\lambda}$, this lemma enables us to make full use of $\mathcal{T}_{\varepsilon,\lambda}$ to treat random Brouwer fixed point theorem. The second tool is the randomized version of the classical Bolzano-Weierstrass theorem [31], which states that an almost surely (briefly, a.s.) bounded sequence in $L^{0}(\mathcal{F},\mathbb{R}^{d})$ always admits an a.s. convergent randomized subsequence although it may admit no a.s. convergent subsequence. Since an a.s. bounded, $\mathcal{T}_{\varepsilon,\lambda}$-closed and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$, is generally noncompact but random sequentially compact by the randomized Bolzano-Weierstrass theorem, then the randomized Bolzano-Weierstrass theorem can be used to overcome the difficulty due to noncompactness [11, 26, 28]. The third tool is a random version of the classical Sperner lemma. To prove a special case of the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Brouwer fixed point theorem (see [11, Proposition 3.1]), the authors [11] first give a special case of random Sperner lemma (see the first part of [11, Theorem 2.3]), where the authors studied a special $L^{0}$-simplicial subdivision of an $L^{0}$-simplex (called the $L^{0}$-barycentric subdivision) in detail. The special random Sperner lemma provided in [11] was employed in all the works [11, 26, 28] on random Brouwer fixed point theorem. However, the proof of [11] on the special random Sperner lemma is not easily accessible to the reader, for example, the idea of the proof in [11] was very similar to Kuhn’s one [35], but the method in [11] was not by means of the graph-theoretic argumentation and thus rather complicated. Considering the fundamental importance of random Brouwer fixed point theorem, for example, it has been used in [26] for the proof of a noncompact Schauder fixed point theorem in random normed modules, and it will also be used to establish a noncompact $KKM$ principle in random locally convex modules so that a series of open problems posed in [26] will be solved in our forthcoming work, it is necessary to look for a very easily accessible proof of random Sperner lemma and to further give a complete proof of random Brouwer fixed point theorem. This leads us directly to systematically developing the theory of $L^{0}$-simplicial subdivisions of $L^{0}$-simplexes.
@@ 2.1 Different forms and special cases of random Brouwer fixed point theorem
It is obvious that a $\mathcal{T}_{\varepsilon,\lambda}$-continuous mapping defined on a $\sigma$-stable set is random sequentially continuous, Proposition 2.3 below, which first occurred in [28, Proposition 4.2] is a special case of Proposition 2.2.
@@
We would also like to point out that Proposition 2.2 (namely Theorem 2.1) and Proposition 2.3 are both best possible, which is explained as follows. It is well known that $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is a metrizable linear topological space under $\mathcal{T}_{\varepsilon,\lambda}$ (namely the topology of convergence in probability), there is, naturally, the notion of boundedness in the sense of the linear topology $\mathcal{T}_{\varepsilon,\lambda}$, denoted by the $\mathcal{T}_{\varepsilon,\lambda}$-boundedness, it is easy to see that the a.s. boundedness implies the $\mathcal{T}_{\varepsilon,\lambda}$-boundedness, and we have known from Proposition 16 of [24] that these two kinds of boundedness coincide for a $\sigma$-stable subset. Likewise, the $L^{0}$-convexity implies the usual convexity, and we have also known from Lemma 2.8 of [49] that these two kinds of convexity coincide for a $\mathcal{T}_{\varepsilon,\lambda}$-closed and $\sigma$-stable subset. On the one hand, an a.s. bounded $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset $G$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is generally noncompact, if the hypothesis on the $\sigma$-stability of $f$ in Proposition 2.3 is removed, then Proposition 2.3 does not hold, which comes from a classical result due to Dobrowolski and Marciszewski [10] (see also [26, Section 2.1]). On the other hand, an example provided by Ponosov in [44] shows that Proposition 2.4 (namely Proposition 2.3) does not hold if $G_{U}$ is replaced with a generic $\mathcal{T}_{\varepsilon,\lambda}$-closed, $\mathcal{T}_{\varepsilon,\lambda}$-bounded and convex set $G$ (namely $G$ is not $L^{0}$-convex). Therefore, they are all best possible.
@@ 2.2 A general $L^{0}$-simplicial subdivision and random Sperner lemma
In [11, Definition 2.1], Drapeau et al. introduced the notion of a proper $L^{0}$-labeling function for the $L^{0}$-barycentric subdivision of an $L^{0}$-simplex in $L^{0}(\mathcal{F},\mathbb{R}^{d})$. Definition 2.7 below is a generalization of [11, Definition 2.1]. But we adopt a strengthened notion of a completely labeled $L^{0}$-simplex since such a notion is both easy to understand and such a completely labeled $L^{0}$-simplex indeed exists (see Theorem 2.9). In fact, according to Definition 2.7, $\phi(y_{j})$ is an equivalence class of a constant valued random variable for each $j\in\{1,2,\cdots,n\}$ for a completely labeled $L^{0}$-simplex $G=Conv_{L^{0}}(\{y_{1},y_{2},\cdots,y_{n}\})$, and hence our notion of a completely labeled $L^{0}$-simplex is completely similar to the notion of a completely labeled simplex.
@@
We first give the following two Examples — Examples 3.1 and 3.2. The first one shows the $L^{0}$-barycentric subdivision introduced in [11, Definition 1.10] is a special case of our Definition 2.6. The second one shows our Definition 2.6 does not just include the $L^{0}$-barycentric subdivision.
@@
Then, by applying the same method as in [11, Lemma 1.11], one can verify that $\{S_{\pi}:\pi\in G(n)\}$ is an $L^{0}$-simplicial subdivision of $S$, which is called the $L^{0}$-barycentric subdivision of $S$.
@@
Definition 4.1 here is adopted from [19, 20] by following the traditional nomenclature of random metric spaces and random normed spaces (see [47, Chapters 9 and 15]).
@@
(1) In Definition 4.3, we require $n_{k}$ to be an element in $L^{0}(\mathcal{F},\mathbb{N})$, rather than a positive integer-valued random variable as in [26, Definition 2.10]. This choice is made for simplifying written expressions, and it is easy to check that the two formulations are essentially equivalent. (2) In the sense of Definition 4.3, the randomized Bolzano-Weierstrass theorem can be restated as: any a.s. bounded nonempty subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is random relatively sequentially compact. Moreover, as pointed out in [26], if $G$ is $\sigma$-stable, then $G$ is random relatively sequentially compact iff $\bar{G}_{\varepsilon,\lambda}$ (the closure of $G$ under $\mathcal{T}_{\varepsilon,\lambda}$) is random sequentially compact. Consequently, any a.s. bounded $\sigma$-stable nonempty $\mathcal{T}_{\varepsilon,\lambda}$-closed subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ (for example, an $L^{0}$-simplex) must be random sequentially compact. (3) For any $(n-1)$-dimensional $L^{0}$-simplex $S$ of an $RN$ module, as pointed out by [26, Lemma 5.5], there exists an $L^{0}$-affine $\mathcal{T}_{\varepsilon,\lambda}$-isomorphism $T$ from $\Delta_{n-1}$ to $S$, where $\Delta_{n-1}:=Conv_{L^{0}}(\{e_{1},e_{2},\cdots,e_{n}\})$ is the standard $(n-1)$-dimensional $L^{0}$-simplex of $L^{0}(\mathcal{F},\mathbb{R}^{n})$. Since $\Delta_{n-1}$ is random sequentially compact as discussed above, and $T$ is $\sigma$-stable and $\mathcal{T}_{\varepsilon,\lambda}$-continuous, then $S$ is random sequentially compact.
@@
It is clear that a.s. sequential continuity implies $\mathcal{T}_{\varepsilon,\lambda}$-continuity. Moreover, when $G_{1}$ is $\sigma$-stable, the $\mathcal{T}_{\varepsilon,\lambda}$-continuity of a mapping $T$ from $G_{1}$ to $G_{2}$ necessarily implies random sequential continuity of $T$. Furthermore, [26, Lemma 4.3] shows that a $\sigma$-stable mapping $T$ of $G_{1}$ to $G_{2}$ is random sequentially continuous iff it is $\mathcal{T}_{c}$-continuous.
@@
Lemma 4.9 is a generalization of [11, Theorem 2.3], since $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is an $RN$ module, and a $\sigma$-stable a.s. sequentially continuous mapping is, obviously, random sequentially continuous, as noted in Remark 4.6. Since random sequential continuity is much more complicated than a.s. sequential continuity, Lemma 4.4 of [26] has played a key role in the treatment of random sequential continuity in the proof of Lemma 4.9. Besides, we would also like to emphasize another difference between the proofs of Lemma 4.9 and [11, Theorem 2.3]: the sequence $\{S^{m},m\in\mathbb{N}\}$ of $(n-1)$-dimensional $L^{0}$-simplexes, which is constructed in Lemma 4.9, has the nice property that $S^{m+1}\subseteq S^{m}$ for each $m\in\mathbb{N}$, but the sequence $\{S^{m},m\in\mathbb{N}\}$ constructed in [11, Theorem 2.3] does not have this property. The nice property of our sequence will be used to study other problems in the future. Besides, the proof of Lemma 4.9 is simpler and clearer than that of [11, Theorem 2.3] since we only employ a single proper $L^{0}$-labeling function throughout the proof of Lemma 4.9.